# Are Foundational Atomistic Models Reliable for Finite-Temperature Molecular Dynamics?

### v2

@@ 5 Conclusion
There is little doubt that we are entering an era where AI is transforming scientific inquiry, particularly in computational materials science. While this Perspective does not attempt to chart the entire frontier, it addresses a focused, practical question: Can we trust foundational atomistic models for finite-temperature MD simulations? Based on our targeted benchmark, the answer is a cautious “yes”, provided their limitations are carefully considered. Many foundational atomistic models demonstrate remarkable accuracy in predicting phonon spectra and equilibrium properties. However, their behavior under realistic MD conditions can be inconsistent, especially in capturing dynamic phase transitions. While our study focuses on a single material, the observed issues likely reflect broader challenges rooted in training data quality, functional choices, and model generalizability. These problems are not intrinsic flaws of the models themselves, but rather symptoms of a still-maturing ecosystem. Moving forward, we see value in hybrid strategies that combine pretraining with targeted fine-tuning, alongside greater emphasis on benchmarking dynamic performance, improving software integration, and fostering community-driven standards. Ultimately, the goal is not to replace specialized models, but to develop robust, adaptable tools that extend the reach of atomistic simulations, unlocking AI’s full potential in computational materials discovery.
6 Methods
All calculations are performed using the ASE 36, with each MLFF integrated as an ASE calculator to compute energy, forces, and stress. To simulate the temperature-driven phase transition of $\mathrm{PbTiO_{3}}$, a $5\times 5\times 5$ supercell containing 625 atoms is constructed from the ground-state structure. MD simulations are carried out in the $NPT$ ensemble using a Parrinello–Rahman barostat coupled with a Nosé–Hoover thermostat. At each temperature, the simulation runs with a 2 fs timestep for 50,000 steps, totaling 100 ps. The last 50 ps of the trajectory is used to compute the averaged lattice constants. Our tests confirm that the cumulative average of the $c/a$ ratio converges after a 10-ps production trajectory. For the $NVT$ ensemble simulation, the same $5\times 5\times 5$ supercell is used, with lattice parameters fixed at their experimental values measured at room temperature. Langevin dynamics are employed with a 2 fs timestep for 50,000 steps. For performance benchmarking, all computations are performed on a single V100-SXM2-16G GPU, with each data point averaged over three independent runs.
7 Code Availability
@@
8 Acknowledgments

### v3

@@ Conclusion
There is little doubt that we are entering an era where AI is transforming scientific inquiry, particularly in computational materials science. While this Perspective does not attempt to chart the entire frontier, it addresses a focused, practical question: Can we trust foundational atomistic models for finite-temperature MD simulations? Based on our targeted benchmark, the answer is a cautious “yes", provided their limitations are carefully considered. Many foundational atomistic models demonstrate remarkable accuracy in predicting phonon spectra and equilibrium properties. However, their behavior under realistic MD conditions can be inconsistent, especially in capturing dynamic phase transitions. While our study focuses on a single material, the observed issues likely reflect broader challenges rooted in training data quality, functional choices, and model generalizability. These problems are not intrinsic flaws of the models themselves, but rather symptoms of a still-maturing ecosystem. Moving forward, we see value in hybrid strategies that combine pretraining with targeted fine-tuning, alongside greater emphasis on benchmarking dynamic performance, improving software integration, and fostering community-driven standards. Ultimately, the goal is not to replace specialized models, but to develop robust, adaptable tools that extend the reach of atomistic simulations, unlocking AI’s full potential in computational materials discovery.
Code Availability
@@
Biographies
Denan Li Denan Li earned his B.S. in Materials Science and Engineering from Ningbo University in 2024. He is currently a Ph.D. candidate in Physics at Westlake University, working in the Multiscale Materials Modeling Laboratory under the supervision of Prof. Shi Liu. His research focuses on the development and application of machine-learning-assisted multiscale methods to understand the dynamic properties of ferroelectrics, with a particular emphasis on organic–inorganic hybrid systems.
Jiyuan Yang Jiyuan Yang earned his B.S. in Physics from Nanjing Normal University in 2017. He received his Ph.D. in Physics from Zhejiang University in 2025 through a joint program with Westlake University. He is currently a postdoctoral fellow at Westlake University, working in Prof. Shi Liu’s lab. His research focuses on the application of deep potential molecular dynamics to study ferroelectric domain dynamics.
Xiangkai Chen Xiangkai Chen received his Ph.D. in Materials Science and Engineering from the University of Science and Technology Beijing in 2023. He is currently a postdoctoral fellow at Westlake University, working in Prof. Shi Liu’s lab. His research focuses on the study of oxide plasticity using the deep potential molecular dynamics method.
Lintao Yu Lintao Yu earned his B.S. in Chemistry from Northeast Petroleum University in 2023. He is currently a Ph.D. candidate in Physics at Westlake University in Prof. Shi Liu’s lab. His research focuses on the design and construction of high-performance computing infrastructure for large-scale simulations.
Shi Liu Shi Liu received his B.S. from the University of Science and Technology of China. He obtained his Ph.D. from the University of Pennsylvania in 2015. He continued his postdoctoral research at the Carnegie Institution for Science in Washington, D.C., and later at the Army Research Laboratory. In June 2019, he joined the Department of Physics at Westlake University, where he is now a tenured Associate Professor. His research interests include novel ferroelectrics, emergent topological phases in condensed matter physics, and deep-learning-based large-scale modeling of complex systems.
Acknowledgments
@@
Table 1: Summary of key properties of various MLFFs used for PTO-test.
Model Version Training Set (Size) Parameters Energy MAE (meV/atom) Force MAE (meV/Å)
CHGNet[Deng23p1031] 0.3.0 MPtrj (146K) 413K 26 60
GPTFF[Xie24p3525] v2 Atomly (37.8M) 502K 32 71
M3GNet[Chen22p718] MP-2021.2.8-PES MPF-2021.2.8 (62.8K) 228K 18.7 63
MACE[Batatia23parXivv1] MP-0b-medium MPtrj (146K) 4.69M 20 45
ORB[Neumann24parXivv1] orb-v2 MPtrj (146K) + Alexandria (3.1M) 25.2M / /
SevenNet[Park24p4857] 7net-l3i5 MPtrj (146K) 1.17M 8.3 29
UniPero[Wu23p180104] v1 Customized (19K) $\approx$500K 1.75 54
Notes: Some models have been updated since their initial publication. When available, the latest version is used, and the energy and force mean absolute errors (MAEs, if reported) are taken from the latest version if available; otherwise, they are taken from the original references. Abbreviations: MPtrj = Materials Project trajectories; MPF = Materials Project structure relaxations
Table 2: Lattice parameters ($a$ and $c$), tetragonality ($c/a$) of the tetragonal PbTiO3 phase, and the energy difference ($\Delta E$) between the tetragonal and cubic phases predicted by different MLFFs. Experimental and DFT results are also included for comparison.
Model $a~(\mathrm{\AA })$ $c~(\mathrm{\AA })$ $c/a$ $\Delta E$ (eV/f.u.)
CHGNet 3.80 5.01 1.32 0.24
GPTFF 3.79 4.88 1.29 0.25
M3GNet 3.80 4.92 1.30 0.16
MACE 3.84 4.83 1.26 0.19
ORB 3.83 4.87 1.27 0.21
SevenNet 3.84 4.79 1.25 0.21
MACE-FT 3.89 4.16 1.07 0.08
UniPero 3.88 4.21 1.08 0.09
LDA[Meyer02p104111] 3.86 4.04 1.05
PBE 3.85 4.73 1.23 0.20
PBEsol 3.87 4.20 1.08 0.08
Exp.[Mabud79p49] 3.90 4.15 1.06
Figure 1: Lattice parameter $a$ and tetragonality ($c/a$) of ground-state PbTiO3 predicted by various MLFFs and exchange-correlation functionals.
Figure 2: Phonon spectra of PbTiO3 calculated using various MLFFs, each based on the optimized ground-state tetragonal structure. The panels show results for: (a) PBE, (b) CHGNet, (c) GPTFF, (d) M3GNet, (e) MACE, (f) ORB, (g) SevenNet, (h) PBEsol, (i) UniPero, and (j) MACE-FT. The spectra obtained from (a) PBE and (h) PBEsol are also included for comparison. (i) UniPero and (j) MACE-FT are trained on a PBEsol-derived database.
Figure 3: Temperature-dependent lattice constants ($a$ and $c$) obtained from NPTNPT MD simulations using various machine learning force fields. The panels show results for: (a) CHGNet, (b) M3GNet, (c) MACE, (d) ORB, (e) SevenNet, (f) GPTFF, (g) MACE-FT, and (h) UniPero. The dashed lines indicate the ground-state lattice parameters ($a_{0}$ and $c_{0}$) of tetragonal PbTiO3 for each force field. The error bars represent the standard deviation over the 50-ps production trajectory, reflecting the extent of thermal fluctuations.
Figure 4: Temperature-dependent spontaneous polarization along the $c$-axis ($P_{z}$) obtained from NVTNVT MD simulations, with lattice parameters fixed to the experimental room-temperature values ($a=3.90$ Å, $c=4.15$ Å). At high temperatures, the polarization does not fully converge to zero due to the imposed tetragonality constraint ($c/a=1.06$). The error bars represent the extent of polarization fluctuations arising from thermal effects over the 50-ps production trajectory.
Figure 5: Computational efficiency benchmark. The reported speed data are for reference only, as a model’s optimal performance can be further improved through careful tuning and the implementation of multi-GPU parallelism.
TOC Graphic